Suppose we had found that the variable giving the amount of grant money raised by the charity in the last year follows a righ please explain related concept in detail for better understansing.
Answer #1

Solution

'According to the Central Limit Theorem, if population is very large, if we perform random sampling and if the the size of the sample is equal to or greater than 30 (n> 30),  regardless of the shape and distribution of the population, the sampling distribution of sample mean will be normal. The value of the sample mean will be approximately equal to the Population mean, and standard deviation of the sample will be : (population std. dev.)/sqrt(n).

a.)Shape: Normal Distribution

b.)Mean(x̅)=μ=500,000

Standard Error=population Std. Dev/sqrt(n)

=850000/sqrt(40)

=134396.8

c.)P(x̅>490000)=1-P(x̅<490000)

=1-P(Z < (490000-500000)/134396.8)

=1-P(Z < -0.07)

=1-0.4721 (use excel function: =NORMSDIST(-0.07))

=0.5279